3x-2y>2 would I subtract 3x or add 3x
Found 2 solutions by user_dude2008, muderman189:
Answer by user_dude2008(1862)

3x-2y>2

-2y>-3x+2

y<(-3x+2)/(-2)

y<(3/2)x-1

Answer by muderman189(2)

you would > turn that to = and do this
Simplifying
3x + -2y = 2
Solving
3x + -2y = 2
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Add '2y' to each side of the equation.
3x + -2y + 2y = 2 + 2y
Combine like terms: -2y + 2y = 0
3x + 0 = 2 + 2y
3x = 2 + 2y
Divide each side by '3'.
x = 0.6666666667 + 0.6666666667y
Simplifying
x = 0.6666666667 + 0.6666666667y this is your answer